determine the general/particular solution for each equation using the applicable solution to equations of order one (separable, homogenous,linear,exact)


  1. (1+y^2)dx + (1=x^2)dy = 0 ; when x = 0, y = 1
1
Expert's answer
2022-02-14T10:46:39-0500

Let us determine the general/particular solution of the differential equation


"(1+y^2)dx + (1+x^2)dy = 0 ;" when "x = 0, y = 1."


After dividing both parts by "(1+y^2) (1+x^2)" we get the equation


"\frac{dx}{1+x^2} + \frac{dy}{1+y^2} = 0,"


and hence


"\int\frac{dx}{1+x^2} + \int\frac{dy}{1+y^2} = C."


It follows that the general solution is of the form:


"\arctan x+\arctan y=C."


Since when "x = 0, y = 1," we get "\arctan 0+\arctan 1=C," and hence "C=\frac{\pi}4."


We conclude that the particular solution is the following:


"\arctan x+\arctan y=\frac{\pi}4."